OpenStudy (anonymous):

Find the sum of the series sum from n=0 to infinity of (-1)^n(8^n)(x^8n)/n!

OpenStudy (anonymous):

Maybe I'm missing something, but you need an actual sum? I'm not sure how you would get a value given you have that x term in there. This isn't a radius/interval of convergence problem or anything?

OpenStudy (anonymous):

I need the actual sum, yes. I think it's asking to simplify (?) it in terms of cosx or sinx or something like that.

OpenStudy (anonymous):

Even if it were a series that represents a known function, without x-values, we just have a function with coefficients determined by n. Unless you mean to say that you think the "sum" isn't a value but a function.

OpenStudy (anonymous):

"Find the sum of the series." Is the actual phrasing. And the question is: \[\sum_{n=0}^{\infty}(-1)^n((8^n)x^(8n))/n!\]

OpenStudy (anonymous):

Hmm, interesting @perl Any idea how this might actually converge to a sum despite the x term?

OpenStudy (anonymous):

I thought I had to manipulate cosx to match it to that, but I can't seem to find a way. :(

OpenStudy (perl):

i take that back

OpenStudy (perl):

is that sum from n=0 to infinity of (-1)^n(8^n)(x^[8n])/n!

OpenStudy (perl):

or

OpenStudy (perl):

sum from n=0 to infinity of (-1)^n(8^n)(x^[8^n])/n!

OpenStudy (perl):

there is a difference

OpenStudy (anonymous):

The latter, but it's x^(8n) ! I'm sorry about that.

OpenStudy (anonymous):

Oh no, it won't be the latter. Haha, sorry.

OpenStudy (perl):

ok that can be summed

OpenStudy (anonymous):

If it's not a hassle, can I ask how you got that?

OpenStudy (perl):

exp ( -8 x^8)

OpenStudy (perl):

why would it be a hassle ? :)

OpenStudy (perl):

its just time consuming :D

OpenStudy (anonymous):

These take so long :(

OpenStudy (perl):

lets use the series for e^x

OpenStudy (perl):

better yet, make it e^u

OpenStudy (perl):

e^u = 1 + u/1! + u^2/2! + u^3/3! + ...

OpenStudy (anonymous):

Okay.

OpenStudy (perl):

the form of that is sum u^n / n!

OpenStudy (anonymous):

Yes.

OpenStudy (perl):

ok so we almost have the correct form

OpenStudy (perl):